Find the number “ c ” that satisfy the Mean Value Theorem (M.V.T) on the given

intervals.

(a) f(x)= e-x , [0, 2]

(b) f(x)= x/(x+2), (1,


1
Expert's answer
2021-06-17T17:43:05-0400

By mean value theorem,

"\text{ There exists} c\in [a,b].\\
f'(c)=\frac{f(b)-f(a)}{b-a}\\
a)
f=e^{-x}\\
f'=-e^{-x}\\
\text{Then, by MVT}\\
f'(c)=\frac{f(2)-f(0)}{2-0}\\
-e^{-c}=\frac{e^{-2}-1}{2-0}\\
-e^{-c}=-0.43\\
-c=ln(0.43)\\
c=0.84\in [0,2].\\
b)\\
f=\frac{x}{x+2}\\
f'=\frac{2}{(x+2)^2}\\
\text{Then, by MVT}\\
f'(c)=\frac{f(2)-f(1)}{2-1}\\
\frac{2}{(c+2)^2}=\frac{\frac{1}{2}
-\frac{1}{3}}{2-1}\\
\frac{2}{(c+2)^2}=\frac{1}{6}\\
(c+2)^2=12\\
c+2=2\sqrt3\\
c=2\sqrt3-2\\
c=1.46\in (1,2)"